# Paul Wild    Lutz Schröder    Karla Messing    Barbara König    Jonas Forster

### v1

@@ 4 Lévy-Prokhorov Distance
The Lévy-Prokhorov distance provides an alternative to the Kantorovich-Wasserstein distance when it comes to measuring the distance between probability distributions. If $(X,d)$ is pseudometric space, and $\mu,\nu\in\mathcal{D}X$ are (discrete) probability distributions, then we define
@@
where $A^{d}_{\varepsilon}=\{y\in Y\mid\inf_{x\in A}d(x,y)\leq\varepsilon\}$. The definition of the Lévy-Prokhorov distance sometimes includes the mirrored condition $\forall B\subseteq Y.\;\nu(B)\leq\mu(B^{d}_{\varepsilon})+\varepsilon$, but this second clause is redundant and does not actually change the induced pseudometric.
@@
The Lévy-Prokhorov distance admits a Wasserstein representation, which is sometimes also known as the Ky Fan metric. It is typically phrased in terms of pairs of (not necessarily independent) random variables that are distributed according to the given distributions, but this is equivalent to considering all their couplings (cf. Section˜2). The predicate lifting $\lambda$ underlying this representation is given by
@@
The predicate lifting from (2) has independently been used under the name ‘generally’ in work on fuzzy description logics [27], and it admits a number of equivalent representations. Intuitively, all of these representations amount to the statement that the value $\lambda_{X}(f)(\mu)$ is high if the value of $f$ is high with high probability.
@@
The Wasserstein representation above is justified by the fact that the predicate lifting is well-behaved:
@@ 4.1 Duality
Next, we show that the Lévy-Prokhorov distance admits a dual representation using the same predicate lifting $\lambda$, that is, we have $K_{\lambda}=W_{\lambda}$. We prove this duality in the more general setting where the two constructions apply to fuzzy relations that need not be pseudometrics. Recall that a fuzzy relation $r\colon X\mathbin{\vtop{\halign{#\cr$\rightarrow$\cr$\hskip 0.51663pt+$\cr}}}Y$ between sets $X$ and $Y$ is a function $r\colon X\times Y\to[0,1]$. The Wasserstein construction applies to fuzzy relations the same way it does to pseudometrics, while the Kantorovich construction is defined in terms of pairs of functions that satisfy a nonexpansiveness condition with respect to the given fuzzy relation. They are therefore both examples of (fuzzy) relational liftings or relators (e.g. [14] and references therein), as they lift fuzzy relations of type $X\mathbin{\vtop{\halign{#\cr$\rightarrow$\cr$\hskip 0.51663pt+$\cr}}}Y$ to relations of type $FX\mathbin{\vtop{\halign{#\cr$\rightarrow$\cr$\hskip 0.51663pt+$\cr}}}FY$:
@@
    The relational Wasserstein lifting $W_{\lambda}(r)\colon FX\mathbin{\vtop{\halign{#\cr$\rightarrow$\cr$\hskip 0.51663pt+$\cr}}}FY$ is defined as
@@
    The relational Kantorovich lifting $K^{\mathsf{rel}}_{\lambda}\colon FX\mathbin{\vtop{\halign{#\cr$\rightarrow$\cr$\hskip 0.51663pt+$\cr}}}FY$ is defined as
@@
Out of the two representations of the Lévy-Prokhorov distance discussed earlier, the second, being a Wasserstein lifting, readily generalizes to fuzzy relations. Therefore, we define the relational Lévy-Prokhorov lifting $\delta^{\mathsf{LP}}$ to be the assignment that maps each fuzzy relation $r\colon X\mathbin{\vtop{\halign{#\cr$\rightarrow$\cr$\hskip 0.51663pt+$\cr}}}Y$ to $\delta^{\mathsf{LP}}(r)=W_{\lambda}(r)\colon\mathcal{D}X\mathbin{\vtop{\halign{#\cr$\rightarrow$\cr$\hskip 0.51663pt+$\cr}}}\mathcal{D}Y$.
We discussed in Section˜3 that the inequality $\leq$’ follows from the general theory of Kantorovich and Wasserstein liftings. The same is true for the respective lax extensions [31, Lemma 5.22], so that it suffices to prove the converse inequality $\geq$’. In the proof of the classical Kantorovich-Rubinstein duality (e.g [30, Theorem 5.10]), this direction amounts to, given an optimal transport plan in the shape of a coupling between the distributions at hand, constructing two price functions that correspond to the optimal cost, in the sense that they form a nonexpansive pair that witnesses the supremum in the definition of the relational Kantorovich lifting $K_{\mathbb{E}}$. In our proof of Lévy-Prokhorov duality we use a similar approach, which means that we should first understand how to phrase computation of the Lévy-Prokhorov distance in terms of a transport problem.
Let $r\colon X\mathbin{\vtop{\halign{#\cr$\rightarrow$\cr$\hskip 0.51663pt+$\cr}}}Y$ and let $\mu\in\mathcal{D}X$ and $\nu\in\mathcal{D}Y$. The Wasserstein representation $\delta^{\mathsf{LP}}(r)(\mu,\nu)=W_{\lambda}(r)(\mu,\nu)$ can be rewritten for this purpose. For $\varepsilon\geq 0$, define $r^{\varepsilon}(x,y)=0$ if $r(x,y)<\varepsilon$ and $r^{\varepsilon}(x,y)=1$ otherwise. Then $\rho(\{(x,y)\mid r(x,y)\geq\varepsilon\})=\mathbb{E}_{\rho}({r^{\varepsilon}})$ for every $\rho\in\mathcal{D}(X\times Y)$, and therefore we have
$W_{\lambda}(r)(\mu,\nu)=\inf\{\inf_{\varepsilon\geq 0}\max(\varepsilon,\mathbb{E}_{\rho}({r^{\varepsilon}}))\mid\rho\in\Gamma(\mu,\nu)\}=\inf_{\varepsilon\geq 0}\max(\varepsilon,W_{\mathbb{E}}(r^{\varepsilon})(\mu,\nu)),$
where we used the representation of $\lambda$ from LABEL:lem:lp-pred-lifting.1 while replacing the strict inequality by a non-strict one and then swapped the infimum over couplings all the way to the inside.
@@
As mentioned before, we only need to prove $W_{\lambda}(r)\leq K^{\mathsf{rel}}_{\lambda}(r)$. Assume $\varepsilon<W_{\lambda}(r)(\mu,\nu)$. Then there is an $r^{\varepsilon}$-nonexpansive pair $(p,q)$ of crisp price functions witnessing the transport cost wrt. $r^{\varepsilon}$. One then replaces the function values 0 and $1$ by $\mathbb{E}_{\mu}({p})$ and $\mathbb{E}_{\mu}({p})+\varepsilon$ to arrive at a $d$-nonexpansive pair $(f,g)$. Using this pair we show
$K^{\mathsf{rel}}_{\lambda}(r)(\mu,\nu)\geq\lambda_{Y}(g)(\nu)-\lambda_{X}(f)(\mu)\geq\mathbb{E}_{\nu}({q})-\mathbb{E}_{\mu}({p})\geq\varepsilon.$
@@ 5 Convex Powerset Functor
A straightforward – but futile – approach to prove duality for $\mathcal{C}$ would be to observe that it holds for the powerset and distribution functor, and then apply a compositionality result. However, the studied liftings (Kantorovich and Wasserstein lifting) are quite fragile when it comes to compositionality, i.e. it does not hold in general that the composition of liftings of functors $F$, $G$ based on the predicate liftings $\lambda^{F}$, $\lambda^{G}$ is the lifting of the composite $FG$ (based on the obvious combined modality $\lambda^{FG}_{X}=\lambda^{F}_{GX}\circ\lambda^{G}_{X}$) [4]. While it is known that the Wasserstein lifting of the convex powerset functor arises by combining the Wasserstein liftings of the component functors [32], this is incorrect for the Kantorovich lifting [9]. In fact, the given counterexample uses a set that is not convex, thus suggesting that the problem might disappear if we restrict to convex sets.
@@
As before, $K_{\lambda}\leq W_{\lambda}$ holds in categorical generality, so the main task is to prove the converse inequality $W_{\lambda}\leq K_{\lambda}$. As discussed earlier, it is known that the Wasserstein representation in terms of $\lambda$ decomposes into the Wasserstein representations in terms of $\sup$ and $\mathbb{E}$, i.e. the Hausdorff and (probabilistic) Kantorovich liftings, respectively:
$W_{\lambda}(d)=W_{\sup}(W_{\mathbb{E}}(d))=\delta^{\mathsf{H}}(\delta^{\mathsf{K}}(d))=:\delta^{\mathsf{HK}}(d)$
@@
Using the above lemma, we may therefore from now on assume that we are working over a metric space $(X,d)$. We may also assume that $X\neq\emptyset$, as otherwise $\mathcal{C}X=\emptyset$ and both Kantorovich and Wasserstein distance are the empty metric, hence equal.
The main intuition behind the proof is best understood in the case where $X=\{x_{1},\dots,x_{n}\}$ is finite, even though the proof will work for arbitrary $X$. In this case we may view probability distributions and fuzzy predicates on $X$ as vectors in n, and the expectation modality simply computes the dot product between two such vectors: $\mathbb{E}_{\mu}({f})=\mu(x_{1})\cdot f(x_{1})+\dots+\mu(x_{n})\cdot f(x_{n}).$ If $A,B\in\mathcal{C}X$ satisfy $\delta^{\mathsf{HK}}(d)(A,B)>\varepsilon$, then this means, by the definition of the Hausdorff distance, that there must be some $\mu\in A$ such that $\delta^{\mathsf{K}}(\mu,\nu)>\varepsilon$ for all $\nu\in B$ (or we are in the symmetric situation with $A$ and $B$ swapped). This implies that $B$ and the $\varepsilon$-ball around $\mu$ are disjoint convex sets, so we can apply the hyperplane separation theorem to find a hyperplane $H$ such that the two sets lie on opposite sides of that plane. A price function witnessing distance at least $\varepsilon$ under the combined modality $\lambda$ can now be constructed from the normal vector of $H$.
@@
In what follows, we often do not distinguish between $\mu$ and $L_{\mu}$ and treat $\mathcal{D}X$ as a subset of $\mathsf{Lip}_{0}(X)^{*}$. We equip $\mathsf{Lip}_{0}(X)^{*}$ with the weak-* topology, which is the weakest topology on $\mathsf{Lip}_{0}(X)^{*}$ making all the maps $\psi\mapsto\psi(f)$ for $f\in\mathsf{Lip}_{0}(X)$ continuous. Equivalently, this is the initial topology wrt. the maps $\psi\mapsto\psi(f)$. Crucially, this topology coincides with the one given by the Kantorovich-Wasserstein distance:
@@
lem:topologies-coincide-discrete Let $(\mu_{n})_{n\in\mathbb{N}}$ be a sequence in $\mathcal{D}X$ and let $\mu\in\mathcal{D}X$. Then $\mu_{n}\to\mu$ in the topology given by $\delta^{\mathsf{K}}(d)$ iff $L_{\mu_{n}}\to L_{\mu}$ in the weak-* topology.
The space $\mathsf{Lip}_{0}(X)^{*}$ is normed via the operator norm $\|\psi\|_{\mathsf{op}}=\sup_{\|f\|_{\mathsf{Lip}}\leq 1}\psi(f)$. This norm relates to the Wasserstein metric as follows:
@@
lem:wasserstein-norm-discrete For any $\mu,\nu\in\mathcal{D}X$ we have $\delta^{\mathsf{K}}(d)(\mu,\nu)=\|\nu-\mu\|_{\mathsf{op}}$.
@@
Let $0<\varepsilon<\delta^{\mathsf{HK}}(d)(A,B)$. As outlined above, we may assume wlog. that there exists some $\mu\in A$ such that $\delta^{\mathsf{K}}(d)(\mu,\nu)>\varepsilon$ for every $\nu\in B$. Let $C$ be the closed $\varepsilon$-ball around $\mu$, shifted by $\mu$, and let $D$ be the closure of $B$, also shifted by $\mu$. These two sets are closed and convex, so by the Hahn-Banach separation theorem there exists a continuous linear functional $g\colon\mathsf{Lip}_{0}(X)^{*}\to\real$ and $c\in\real$ such that
@@
The range of $f_{1}$ is not necessarily contained in $[0,1]$, but it must be contained in some subinterval of of length at most $1$ by nonexpansiveness and because $d$ is $1$-bounded. As expectation is linear, we may simply shift $f_{1}$ by a suitable amount. ∎
@@
It has been shown in work on correspondences between Kantorovich and Wasserstein representations of metric liftings [16] (cf. Section 1 under related work) that such correspondences can be combined along sums and product of functors, so that one arrives at general correspondence results for classes of functors obtained by closing given basic building blocks (originally constant functors, identity, powerset, and distributions) under sum and product. One thus obtains correspondences for composite system types such as labelled Markov chains [16, Example 41]. The correspondences produced in this way are not generalized Kantorovich-Rubinstein dualities in the strict sense we use here, as the transition from the Wasserstein presentation to the Kantorovich presentation in general involves the introduction of additional modalities. In particular, this happens for products, where one needs to introduce separate modalities for the factors (indeed, this is what is morally behind the fact that generalized Kantorovich-Rubinstein duality fails for the squaring functor [4]). Nevertheless, our results on generalized Kantorovich-Rubinstein duality for Lévy-Prokhorov distance and convex powerset imply that these functors can now be used as additional basic building blocks in this framework.

### v2

@@ 4 Lévy-Prokhorov Distance
The Lévy-Prokhorov distance provides an alternative to the Kantorovich-Rubinstein distance when it comes to measuring the distance between probability distributions. If $(X,d)$ is a pseudometric space, and $\mu,\nu\in\mathcal{D}X$ are (discrete) probability distributions, then we define
@@
where $A^{d}_{\varepsilon}=\{y\in X\mid\inf_{x\in A}d(x,y)\leq\varepsilon\}$. The definition of the Lévy-Prokhorov distance sometimes includes the mirrored condition $\forall B\subseteq X.\;\nu(B)\leq\mu(B^{d}_{\varepsilon})+\varepsilon$, but this second clause is redundant and does not actually change the induced pseudometric.
@@
The Lévy-Prokhorov distance admits a representation based on couplings (cf. Section˜3), or equivalently in terms of pairs of (not necessarily independent) random variables that are distributed according to the given distributions, known as the Ky Fan metric [12]. The predicate lifting $\lambda$ underlying this representation is given by
@@
The predicate lifting from (3) has independently been used under the name ‘generally’ in work on fuzzy description logics [34], and it admits a number of equivalent representations. Intuitively, all of these representations amount to the statement that the value $\lambda_{X}(f)(\mu)$ is high if the value of $f$ is high with high probability when sampling according to the distribution $\mu$.
@@
The coupling-based representation above is justified by the fact that the predicate lifting is well-behaved:
@@ 4.1 Duality
Next, we show that the Lévy-Prokhorov distance admits a dual representation using the same predicate lifting $\lambda$, that is, we have $K_{\lambda}=W_{\lambda}$. We prove this duality in the more general setting where the two constructions apply to fuzzy relations that need not be pseudometrics. Recall that a fuzzy relation $r\colon X\mathbin{\vtop{\halign{#\cr$\rightarrow$\cr$\hskip 0.51663pt+$\cr}}}Y$ between sets $X$ and $Y$ is a function $r\colon X\times Y\to[0,1]$. The coupling-based construction applies to fuzzy relations the same way it does to pseudometrics, while the price-function-based construction is defined in terms of pairs of functions that satisfy a nonexpansiveness condition with respect to the given fuzzy relation. They are therefore both examples of (fuzzy) relational liftings or relators (e.g. [15] and references therein), as they lift fuzzy relations of type $X\mathbin{\vtop{\halign{#\cr$\rightarrow$\cr$\hskip 0.51663pt+$\cr}}}Y$ to relations of type $FX\mathbin{\vtop{\halign{#\cr$\rightarrow$\cr$\hskip 0.51663pt+$\cr}}}FY$:
@@
    The relational coupling-based lifting $W_{\lambda}(r)\colon FX\mathbin{\vtop{\halign{#\cr$\rightarrow$\cr$\hskip 0.51663pt+$\cr}}}FY$ is defined as
@@
    The relational price-function-based lifting $K^{\mathsf{rel}}_{\lambda}\colon FX\mathbin{\vtop{\halign{#\cr$\rightarrow$\cr$\hskip 0.51663pt+$\cr}}}FY$ is defined as
@@
Out of the two representations of the Lévy-Prokhorov distance discussed earlier, the second, being based on couplings, readily generalizes to fuzzy relations. Therefore, we define the relational Lévy-Prokhorov lifting $\delta^{\mathsf{LP}}$ to be the assignment that maps each fuzzy relation $r\colon X\mathbin{\vtop{\halign{#\cr$\rightarrow$\cr$\hskip 0.51663pt+$\cr}}}Y$ to $\delta^{\mathsf{LP}}(r)=W_{\lambda}(r)\colon\mathcal{D}X\mathbin{\vtop{\halign{#\cr$\rightarrow$\cr$\hskip 0.51663pt+$\cr}}}\mathcal{D}Y$.
We discussed in Section˜3 that the inequality $\leq$’ follows from the general theory of coupling-based and price-function-based liftings. The same is true for the respective lax extensions [38, Lemma 5.22], so that it suffices to prove the converse inequality $\geq$’. In the proof of the classical Kantorovich-Rubinstein duality (e.g [37, Theorem 5.10]), this direction amounts to, given an optimal transport plan in the shape of a coupling between the distributions at hand, constructing two price functions that correspond to the optimal cost, in the sense that they form a nonexpansive pair that witnesses the supremum in the definition of the relational price-function-based lifting $K_{\mathbb{E}}$. In our proof of Lévy-Prokhorov duality we use a similar approach, which means that we should first understand how to phrase computation of the Lévy-Prokhorov distance in terms of a transport problem.
Let $r\colon X\mathbin{\vtop{\halign{#\cr$\rightarrow$\cr$\hskip 0.51663pt+$\cr}}}Y$ and let $\mu\in\mathcal{D}X$ and $\nu\in\mathcal{D}Y$. The coupling-based representation $\delta^{\mathsf{LP}}(r)(\mu,\nu)=W_{\lambda}(r)(\mu,\nu)$ can be rewritten for this purpose. For $\varepsilon\geq 0$, define $r^{\varepsilon}(x,y)=0$ if $r(x,y)<\varepsilon$ and $r^{\varepsilon}(x,y)=1$ otherwise. Then $\rho(\{(x,y)\mid r(x,y)\geq\varepsilon\})=\mathbb{E}_{\rho}({r^{\varepsilon}})$ for every $\rho\in\mathcal{D}(X\times Y)$. We may now use the representation of $\lambda$ from LABEL:lem:lp-pred-lifting.1, replacing the strict inequality with a non-strict one, and then swap the infimum over couplings inside to obtain
$W_{\lambda}(r)(\mu,\nu)=\inf\{\inf_{\varepsilon\geq 0}\max(\varepsilon,\mathbb{E}_{\rho}({r^{\varepsilon}}))\mid\rho\in\Gamma(\mu,\nu)\}=\inf_{\varepsilon\geq 0}\max(\varepsilon,W_{\mathbb{E}}(r^{\varepsilon})(\mu,\nu)).$
@@
As mentioned before, we only need to prove $W_{\lambda}(r)\leq K^{\mathsf{rel}}_{\lambda}(r)$. Assume $\varepsilon<W_{\lambda}(r)(\mu,\nu)$. Then there is an $r^{\varepsilon}$-nonexpansive pair $(p,q)$ of crisp price functions witnessing the transport cost wrt. $r^{\varepsilon}$. One then replaces the function values 0 and $1$ by $\mathbb{E}_{\mu}({p})$ and $\mathbb{E}_{\mu}({p})+\varepsilon$ to arrive at an $r$-nonexpansive pair $(f,g)$. Using this pair we show
$K^{\mathsf{rel}}_{\lambda}(r)(\mu,\nu)\geq\lambda_{Y}(g)(\nu)-\lambda_{X}(f)(\mu)\geq(\mathbb{E}_{\mu}({p})+\varepsilon)-\mathbb{E}_{\mu}({p})=\varepsilon.$
@@ 5 Convex Powerset Functor
A straightforward – but futile – approach to prove duality for $\mathcal{C}$ would be to observe that it holds for the powerset and distribution functor, and then apply a compositionality result. However, the studied liftings (price-function-based and coupling-based) are quite fragile when it comes to compositionality, i.e. it does not hold in general that the composition of liftings of functors $F$, $G$ based on the predicate liftings $\lambda^{F}$, $\lambda^{G}$ is the lifting of the composite $FG$ (based on the obvious combined modality $\lambda^{FG}_{X}=\lambda^{F}_{GX}\circ\lambda^{G}_{X}$) [4]. While it is known that the coupling-based lifting of the convex powerset functor arises by combining the coupling-based liftings of the component functors [39], this is incorrect for the price-function-based lifting [9]. In fact, the given counterexample uses a set that is not convex, thus suggesting that the problem might disappear if we restrict to convex sets.
@@
As before, $K_{\lambda}\leq W_{\lambda}$ holds in categorical generality, so the main task is to prove the converse inequality $W_{\lambda}\leq K_{\lambda}$. As discussed earlier, it is known that the coupling-based representation in terms of $\lambda$ decomposes into the coupling-based representations in terms of $\sup$ and $\mathbb{E}$, i.e. the Hausdorff and Kantorovich-Rubinstein liftings, respectively:
$W_{\lambda}(d)=W_{\sup}(W_{\mathbb{E}}(d))=\delta^{\mathsf{H}}(\delta^{\mathsf{KR}}(d))=:\delta^{\mathsf{HK}}(d)$
@@
Using the above lemma, we may therefore from now on assume that we are working over a metric space $(X,d)$. We may also assume that $X\neq\emptyset$, as otherwise $\mathcal{C}X=\emptyset$ and both coupling-based and price-function-based distance are the empty metric, hence equal.
The main intuition behind the proof is best understood in the case where $X=\{x_{1},\dots,x_{n}\}$ is finite, even though the proof will work for arbitrary $X$. In this case we may view probability distributions and fuzzy predicates on $X$ as vectors in n, and the expectation modality simply computes the dot product between two such vectors: $\mathbb{E}_{\mu}({f})=\mu(x_{1})\cdot f(x_{1})+\dots+\mu(x_{n})\cdot f(x_{n}).$ If $A,B\in\mathcal{C}X$ satisfy $\delta^{\mathsf{HK}}(d)(A,B)>\varepsilon$, then this means, by the definition of the Hausdorff distance, that there must be some $\mu\in A$ such that $\delta^{\mathsf{KR}}(\mu,\nu)>\varepsilon$ for all $\nu\in B$ (or we are in the symmetric situation with $A$ and $B$ swapped). This implies that $B$ and the $\varepsilon$-ball around $\mu$ are disjoint convex sets, so we can apply the hyperplane separation theorem to find a hyperplane $H$ such that the two sets lie on opposite sides of that plane (more precisely, $B$ may intersect the hyperplane, but the $\varepsilon$-ball may not). A price function witnessing distance at least $\varepsilon$ under the combined modality $\lambda$ can now be constructed from the normal vector of $H$.
@@
In what follows, we often do not distinguish between $\mu$ and $L_{\mu}$ and treat $\mathcal{D}X$ as a subset of $\mathsf{Lip}_{0}(X)^{*}$. We equip $\mathsf{Lip}_{0}(X)^{*}$ with the weak-* topology, which is the weakest topology on $\mathsf{Lip}_{0}(X)^{*}$ making all the maps $\psi\mapsto\psi(f)$ for $f\in\mathsf{Lip}_{0}(X)$ continuous. Equivalently, this is the initial topology wrt. the maps $\psi\mapsto\psi(f)$. Crucially, this topology coincides with the one given by the Kantorovich-Rubinstein distance:
@@
lem:topologies-coincide-discrete Let $(\mu_{n})_{n\in\mathbb{N}}$ be a sequence in $\mathcal{D}X$ and let $\mu\in\mathcal{D}X$. Then $\mu_{n}\to\mu$ in the topology given by $\delta^{\mathsf{KR}}(d)$ iff $L_{\mu_{n}}\to L_{\mu}$ in the weak-* topology.
The space $\mathsf{Lip}_{0}(X)^{*}$ is normed via the operator norm $\|\psi\|_{\mathsf{op}}=\sup_{\|f\|_{\mathsf{Lip}}\leq 1}\psi(f)$. This norm relates to the Kantorovich-Rubinstein metric as follows:
@@
lem:wasserstein-norm-discrete For any $\mu,\nu\in\mathcal{D}X$ we have $\delta^{\mathsf{KR}}(d)(\mu,\nu)=\|\nu-\mu\|_{\mathsf{op}}$.
@@
Let $0<\varepsilon<\delta^{\mathsf{HK}}(d)(A,B)$. As outlined above, we may assume wlog. that there exists some $\mu\in A$ such that $\delta^{\mathsf{KR}}(d)(\mu,\nu)>\varepsilon$ for every $\nu\in B$. Let $C$ be the closed $\varepsilon$-ball around $\mu$, shifted by $\mu$, and let $D$ be the closure of $B$, also shifted by $\mu$. These two sets are closed and convex, so by the Hahn-Banach separation theorem there exists a continuous linear functional $g\colon\mathsf{Lip}_{0}(X)^{*}\to\real$ and $c\in\real$ such that
@@
The range of $f_{1}$ is not necessarily contained in $[0,1]$, but it must be contained in some subinterval of of length at most $1$ by nonexpansiveness and because $d$ is $1$-bounded. As expectation is linear, we may simply shift $f_{1}$ by a suitable amount to extract the desired price function. ∎
@@
It has been shown in work on correspondences between price-function-based and coupling-based representations of metric liftings [23] (cf. Section 1 under related work) that such correspondences can be combined along sums and product of functors, so that one arrives at general correspondence results for classes of functors obtained by closing given basic building blocks (originally constant functors, identity, powerset, and distributions) under sum and product. One thus obtains correspondences for composite system types such as labelled Markov chains [23, Example 41]. The correspondences produced in this way are not generalized Kantorovich-Rubinstein dualities in the strict sense we use here, as the transition from the coupling-based presentation to the price-function-based presentation in general involves the introduction of additional modalities. In particular, this happens for products, where one needs to introduce separate modalities for the factors (indeed, this is what is morally behind the fact that generalized Kantorovich-Rubinstein duality fails for the squaring functor [4]). Nevertheless, our results on generalized Kantorovich-Rubinstein duality for Lévy-Prokhorov distance and convex powerset imply that these functors can now be used as additional basic building blocks in this framework.